Let X be a random number from (0,1). Find the probability density functions of the random variables Y = -ln(1-X) and Z =X^{n}, n\neq 0
Answer #1

X be random number from (0,1) . So it can be considered as uniform. We can write the density as

f_X(x)=1\quad 0<x<1.

Let Y=-\ln (1-X)

Y=-\ln (1-X)\\ \Rightarrow -Y=\ln(1-X)\\ \Rightarrow e^{-Y}=1-X\\ or \quad X=1-e^{-Y}

When x=0, y=0 and when x=1 , y=\infty

\frac{dx}{dy}=e^{-y}

So the density of Y is given by

f_Y(y)=f_X(x)\left|\frac{dx}{dy}\right|\\=1.e^{-y}\\ =e^{-y},\quad 0<y<\inftywhich is the exponential density.

So Y=-\ln(1-X) has exponential distribution with probability density function

f_Y(y)=e^{-y},\quad 0<y<\infty.

Now we want to find the density of Z=X^n

Whenx=0,z=0 and when x=1, z=1.

Z=X^n\Rightarrow X=Z^{\frac{1}{n}}

\frac{dx}{dz}=\frac{1}{n}z^{\frac{1}{n}-1}=\frac{1}{n}z^{\frac{1-n}{n}}

So the density of Z is given by

f_Z(z)=f_X(x)\left|\frac{dx}{dz}\right|\\ =1.\frac{1}{n}z^{\frac{1-n}{n}}\\ =\frac{1}{n}z^{\frac{1-n}{n}},\quad 0<z<1

So the probability density function of Z is given by f_Z(z)=\frac{1}{n}z^{\frac{1-n}{n}},\quad 0<z<1 .